# Supporting the development of Machine Learning for fundamental science in a federated Cloud with the AI_INFN platform

### v1

@@ Abstract
Machine Learning (ML) is driving a revolution in the way scientists design, develop, and deploy data-intensive software. However, the adoption of ML presents new challenges for the computing infrastructure, particularly in terms of provisioning and orchestrating access to hardware accelerators for development, testing, and production. The INFN-funded project $\mathrm{AI\_INFN}$ (“Artificial Intelligence at INFN”) aims at fostering the adoption of ML techniques within INFN use cases by providing support on multiple aspects, including the provision of AI-tailored computing resources. It leverages cloud-native solutions in the context of INFN Cloud, to share hardware accelerators as effectively as possible, ensuring the diversity of the Instituteâs research activities is not compromised. In this contribution, we provide an update on the commissioning of a $\mathrm{Kubernetes}$ platform designed to ease the development of GPU-powered data analysis workflows and their scalability on heterogeneous, distributed computing resources, possibly federated as Virtual Kubelets with the interLink provider.
@@ 1 Introduction
Cloud computing has also played a key role in accelerating the adoption of AI techniques by making computing power and data storage more accessible to developers by outsourcing resource management to service providers like AWS [^1] Amazon Web Services, https://aws.amazon.com. or GCP [^2] Google Cloud Platform, https://cloud.google.com.. This approach enables ready-to-use software to be delivered directly to the end users, a model known as Software-as-a-Service (SaaS). A typical example is JupyterHub [^3] Multi-user platform for collaborative coding and data analysis using notebooks, https://jupyter.org/hub., which provides multiple users with access to notebooks for data visualization and interactive computation using resources provisioned by the Cloud. The growing community of Data Scientists makes large use of this or similar technologies, contributing to expanding the catalog of libraries and tools available for developing and deploying AI applications.
@@ 2 The AI_INFN initiative
The $\mathrm{AI\_INFN}$ platform is a SaaS provided by $\mathrm{INFN\ Cloud}$. The underlying specialized and dedicated hardware was designed for High Performance Computing (HPC) tasks and is hosted and managed at INFN CNAF in Bologna, few steps away from the pre-exascale CINECA supercomputer Leonardo [^5] CINECA Leonardo, https://leonardo-supercomputer.cineca.eu. The infrastructure comprises four servers, clustered in an OpenStack [^6] OpenStack, https://www.openstack.org. tenancy, acquired and installed between 2020 and 2024 as the demand for Cloud-accessible HPC computing resources increased:
@@
At the time of writing, 72 researchers working on 16 research activities have requested and gained access to the platform. On average, 10 to 15 researchers connect at least once to the platform in a working day. Two dedicated clones of the $\mathrm{AI\_INFN}$ platform were temporarily deployed at CNAF and at ReCaS Bari to provision the GPU-accelerated resources to the 30 participants of the first AI_INFN hackathon, an advanced training event organized in Padua in November 2024.
@@ 3 SaaS provisioning model, the AI_INFN platform
[…]
The main platform file system is distributed through the containers via NFS. One of the platform nodes runs an NFS server in a $\mathrm{Kubernetes}$ pod and exports data to the containers spawned by $\mathrm{JupyterHub}$. At spawn time, $\mathrm{JupyterHub}$ is configured to create the user’s home directories and project-dedicated shared volumes. A special directory of the platform file system, that users can use directly or clone and extend in their directories, is reserved for distributing managed software environments, configure using virtual environments. The platform file system is subject to regular encrypted backup. Backup data is stored in a remote Ceph ceph volume provisioned by $\mathrm{INFN\ Cloud}$ using the BorgBackup [^9] BorgBackup, https://borgbackup.readthedocs.io/en/stable/#. package to ensure data deduplication.
@@
At the opposite extreme of the I/O performance spectrum there are distributed virtual file systems that can be mounted on multiple computing resources, enabling the sharing of notebooks and user-defined computing environments across multiple computing sites and compute backends. JuiceFS [^12] JuiceFS, https://juicefs.com/en. is a cloud-based, high-performance, POSIX-compliant distributed file system specifically designed for multi-cloud and serverless computing. It decouples data and metadata delegating these tasks to highly optimized third party projects, combining a metadata engine implemented with either key-value databases (such as Redis [^13] Redis, https://redis.io.) or relational database management systems (such as PostgreSQL [^14] PostgreSQL, https://www.postgresql.org.) with storage systems accessed through S3, WebDAV or other high-throughput protocols.
@@
Finally users can install or upgrade packages in their containers. Installing new software will introduce ephemeral modifications in OverlayFS layer on top of the container file system. A more effective and popular alternative to installing packages in the container is to rely on the binaries distributed through the CERN VM file system (cvmfs). CVMFS, used to distribute software through the nodes of the WLCG, is made available to the platform users through a $\mathrm{Kubernetes}$ installation that shares the caches among different users and sessions.
@@
In the $\mathrm{AI\_INFN}$ platform, the NVIDIA GPU operator [^15] NVIDIA GPU Operator, https://docs.nvidia.com/datacenter/cloud-native/gpu-operator/latest/index.html., is used to install and maintain the GPU drivers. The NVIDIA GPU Operator on $\mathrm{Kubernetes}$ streamlines the management and deployment of NVIDIA GPU resources by automating the installation and configuration of required components within a $\mathrm{Kubernetes}$ cluster. In general, the GPU Operator is designed to simplify the management of clusters with a large number of GPU accelerators, enabling efficient scaling and resource optimization for GPU-accelerated workloads such as AI, ML, and data processing. This approach enables centralized management of GPUs, ensuring a consistent and scalable configuration across all nodes in the cluster, while simplifying maintenance and updates.
One of the most common support requests with the VM-based provisioning model involved setting up a GPU-accelerated Python software stack. While the TOSCA template TOSCA and Ansible playbook handled the installation of the NVIDIA driver and runtime, choosing and installing the Python libraries required for the application was left to the users.
Typically, managing a data science project’s dependencies is the responsibility of developers and analysts, and many analysis projects require multiple computing environments. When introducing the $\mathrm{AI\_INFN}$ platform, particular attention was given to ensuring user sessions were highly customizable and adaptable by providing mechanisms for users to create and manage their own computing environments. The most radical customization option is to build and pick a custom OCI image. Both communities and individual users can modify the default OCI image by adding system libraries or software packages or by altering the $\mathrm{JupyterLab}$ service itself.
While users often prefer conda [^16] Conda, https://conda.io. for custom software environments, Apptainer [^17] Apptainer, https://apptainer.org/. images are gaining popularity. Unlike conda, which consists of thousands of small files, Apptainer uses SquashFS [^18] Squashfs, https://www.kernel.org/doc/Documentation/filesystems/squashfs.txt., a compressed read-only file system, to package the entire environment into a single file. This makes Apptainer images easier to share and distribute through object stores.
The $\mathrm{AI\_INFN}$ platform provides documentation to help users export conda environments as Apptainer images and use them as Jupyter kernels. It also offers pre-built conda environments and Apptainer images with software versions optimized for GPU-accelerated Machine Learning frameworks. Users can clone these environments and add project-specific dependencies, typically related to data loading and visualization, and independent of the GPU software stack.
A notable exception is represented by the software environment to develop Quantum Machine Learning (QML), featuring Python modules that simulate the effect of quantum operators on GPU and therefore requiring the same attention as other GPU-accelerated ML libraries to match the versions of the underlying software.
In addition, Apptainer images specialized for the data processing of the LHC experiments can be obtained via CVMFS.
A dedicated monitoring and accounting system has been set up for the platform in order to effectively control the use of all the platform’s resources and in particular of the GPUs.
@@
Several metric exporters have been configured to collect the information of interest and then expose it to a Prometheus [^19] Prometheus, https://prometheus.io instance running in the platform. Some of these exporters are already available as Free and Open Source Software, such as Kube Eagle [^20] Martin Schneppenheim, Kube Eagle (2020). GitHub repository: https://github.com/cloudworkz/kube-eagle., which manages information about the use of the cluster’s CPUs and memory resources by the various components of $\mathrm{Kubernetes}$, or NVIDIA GPU DCGM exporter [^21] NVIDIA DCGM Exporter, https://docs.nvidia.com/datacenter/cloud-native/gpu-telemetry/latest/dcgm-exporter.html.. Other exporters were developed on purpose, for example to monitor the usage of storage resources. All the metrics collected by Prometheus are then made visible and accessible through a Grafana [^22] Grafana Labs: Grafana Documentation, https://grafana.com/docs/. dashboard. Grafana is run in a VM independent of the platform cluster and is used to monitor other VMs in the $\mathrm{AI\_INFN}$ OpenStack tenancy. It also hosts a PostgreSQL database for the accounting metrics, updated at regular intervals by averaging the metrics obtained from the monitoring Prometheus service.
@@ 4 Offloading: scale the applications beyond cluster boundaries
Users are allowed to scale beyond their notebook instance by creating $\mathrm{Kubernetes}$ jobs, enqueued and assigned to either local or remote resources by the Kueue controller [^23] Kueue, https://kueue.sigs.k8s.io/.. Kueue is designed to use local resources in an opportunistic way, configuring the running batch jobs to be immediately evicted in case new notebook instances are spawned pushing the cluster in a condition of resource contention. User do not create jobs directly accessing $\mathrm{Kubernetes}$ APIs, but passing through a dedicated microservice, named vkd, that validates user’s request based on memberships criteria and manage $\mathrm{Kubernetes}$ secrets that are not intended to be exposed to users, but still are needed for their jobs to be executed in the platform.
@@
Virtual nodes are $\mathrm{Kubernetes}$ nodes that are not backed by a Linux kernel but mimic a $\mathrm{Kubernetes}$ kubelet in the interactions with the $\mathrm{Kubernetes}$ API server. The software component providing this interface is named Virtual Kubelet [^24] Virtual Kubelet, https://virtual-kubelet.io/. and is designed to ease the integration with various resource providers. The $\mathrm{AI\_INFN}$ platform relies on the InterLink Ciangottini:chep2024 provider. A further abstraction layer defining a simplified set of REST APIs that can be implemented by the so-called InterLink plugins providing the actual access to the compute resources. At the time of writing, the $\mathrm{AI\_INFN}$ platform is interfaced with plugins accessing HTCondor htcondor , Slurm slurm and Podman [^25] Podman, https://podman.io/. resources. Following a recent integration test, a $\mathrm{Kubernetes}$ plugin will be brought to production soon. A schematic representation of the architecture is provided in Figure 2.
@@
Figure 2 reports a recent scalability test involving resources provisioned by four different sites, without distributing the file system and for CPU-only payloads of the LHCb Flash Simulation Barbetti:2906203 .
Figure 1: Enabling offloading using interLink as Virtual Kubelet provider.

### v2

@@ Abstract
Machine Learning (ML) is driving a revolution in the way scientists design, develop, and deploy data-intensive software. However, the adoption of ML presents new challenges for the computing infrastructure, particularly in terms of provisioning and orchestrating access to hardware accelerators for development, testing, and production. The INFN-funded project $\mathrm{AI\_INFN}$ (“Artificial Intelligence at INFN”) aims at fostering the adoption of ML techniques within INFN use cases by providing support on multiple aspects, including the provision of AI-tailored computing resources. It leverages cloud-native solutions in the context of INFN Cloud, to share hardware accelerators as effectively as possible, ensuring the diversity of the Institute’s research activities is not compromised. In this contribution, we provide an update on the commissioning of a $\mathrm{Kubernetes}$ platform designed to ease the development of GPU-powered data analysis workflows and their scalability on heterogeneous, distributed computing resources, possibly federated as Virtual Kubelets with the interLink provider.
@@ 1 Introduction
Cloud computing has also played a key role in accelerating the adoption of AI techniques by making computing power and data storage more accessible to developers by outsourcing resource management to service providers like AWS AWS or GCP GCP . This approach enables ready-to-use software to be delivered directly to the end users, a model known as Software-as-a-Service (SaaS). A typical example is JupyterHub JupyterHub , which provides multiple users with access to notebooks for data visualization and interactive computation using resources provisioned by the Cloud. The growing community of Data Scientists makes large use of this or similar technologies, contributing to expanding the catalog of libraries and tools available for developing and deploying AI applications.
@@ 2 The AI_INFN initiative
The $\mathrm{AI\_INFN}$ platform is a SaaS provided by $\mathrm{INFN\ Cloud}$. The underlying specialized and dedicated hardware was designed for High Performance Computing (HPC) tasks and is hosted and managed at INFN CNAF in Bologna, few steps away from the pre-exascale CINECA supercomputer Leonardo [^2] More details available at https://leonardo-supercomputer.cineca.eu. The infrastructure comprises four servers, clustered in an OpenStack OpenStack tenancy, acquired and installed between 2020 and 2024 as the demand for Cloud-accessible HPC computing resources increased:
@@
At the time of writing, 72 researchers working on 16 research activities have requested and gained access to the platform. On average, 10 to 15 researchers connect at least once to the platform in a working day. Two dedicated clones of the $\mathrm{AI\_INFN}$ platform were temporarily deployed at CNAF and at ReCaS Bari to provision the GPU-accelerated resources to the 30 participants of the first AI_INFN hackathon, an advanced training event organized in Padua in November 2024333More details available at https://agenda.infn.it/event/43129..
@@ 3 SaaS provisioning model, the AI_INFN platform
The $\mathrm{AI\_INFN}$ platform was deployed on a $\mathrm{Kubernetes}$ Kubernetes cluster spanning on at least three VMs within the dedicated OpenStack tenancy providing part of the storage resource, the monitoring infrastructure and the $\mathrm{Kubernetes}$ control plane. A minimal amount of compute resources is also provisioned to make it possible for users to access their data on the platform anytime Anderlini:wmlq2024 . Additional compute resource provided by VMs can be attached to the cluster and detached to be used as standalone machines running an Ansible Ansible playbook, or reassigned to another cluster in the same tenancy. $\mathrm{AI\_INFN}$ users are identified through INFN Cloud Indigo IAM indigoiam instance. Once authenticated, users can configure and spawn their $\mathrm{JupyterLab}$ instance using $\mathrm{JupyterHub}$.
The main platform file system is distributed through the containers via NFS. One of the platform nodes runs an NFS server in a $\mathrm{Kubernetes}$ pod and exports data to the containers spawned by $\mathrm{JupyterHub}$. At spawn time, $\mathrm{JupyterHub}$ is configured to create the user’s home directories and project-dedicated shared volumes. A special directory of the platform file system, that users can use directly or clone and extend in their directories, is reserved for distributing managed software environments, configure using virtual environments. The platform file system is subject to regular encrypted backup. Backup data is stored in a remote Ceph ceph volume provisioned by $\mathrm{INFN\ Cloud}$ using the BorgBackup BorgBackup package to ensure data deduplication.
@@
At the opposite extreme of the I/O performance spectrum there are distributed virtual file systems that can be mounted on multiple computing resources, enabling the sharing of notebooks and user-defined computing environments across multiple computing sites and compute backends. JuiceFS JuiceFS is a cloud-based, high-performance, POSIX-compliant distributed file system specifically designed for multi-cloud and serverless computing. It decouples data and metadata delegating these tasks to highly optimized third party projects, combining a metadata engine implemented with either key-value databases (such as Redis redis-link ) or relational database management systems (such as PostgreSQL PostgreSQL ) with storage systems accessed through S3, WebDAV or other high-throughput protocols.
@@
Finally users can install or upgrade packages in their containers. Installing new software will introduce ephemeral modifications in OverlayFS layer on top of the container file system. A more effective and popular alternative to installing packages in the container is to rely on the binaries distributed through the CERN VM file system (cvmfs) Buncic:2010zz . CVMFS, used to distribute software through the nodes of the WLCG, is made available to the platform users through a $\mathrm{Kubernetes}$ installation that shares the caches among different users and sessions.
@@
In the $\mathrm{AI\_INFN}$ platform, the NVIDIA GPU operator GpuOperator , is used to install and maintain the GPU drivers. The NVIDIA GPU Operator on $\mathrm{Kubernetes}$ streamlines the management and deployment of NVIDIA GPU resources by automating the installation and configuration of required components within a $\mathrm{Kubernetes}$ cluster. In general, the GPU Operator is designed to simplify the management of clusters with a large number of GPU accelerators, enabling efficient scaling and resource optimization for GPU-accelerated workloads such as AI, ML, and data processing. This approach enables centralized management of GPUs, ensuring a consistent and scalable configuration across all nodes in the cluster, while simplifying maintenance and updates.
One of the most common support requests with the VM-based provisioning model involved setting up a GPU-accelerated Python software stack. While the TOSCA template TOSCA and Ansible playbook handled the installation of the NVIDIA driver and runtime, choosing and installing the Python libraries required for the application was left to the users.Typically, managing a data science project’s dependencies is the responsibility of developers and analysts, and many analysis projects require multiple computing environments. When introducing the $\mathrm{AI\_INFN}$ platform, particular attention was given to ensuring user sessions were highly customizable and adaptable by providing mechanisms for users to create and manage their own computing environments. The most radical customization option is to build and pick a custom OCI image. Both communities and individual users can modify the default OCI image by adding system libraries or software packages or by altering the $\mathrm{JupyterLab}$ service itself.While users often prefer conda conda for custom software environments, Apptainer Apptainer images are gaining popularity. Unlike conda, which consists of thousands of small files, Apptainer uses SquashFS Squashfs , a compressed read-only file system, to package the entire environment into a single file. This makes Apptainer images easier to share and distribute through object stores.The $\mathrm{AI\_INFN}$ platform provides documentation to help users export conda environments as Apptainer images and use them as Jupyter kernels. It also offers pre-built conda environments and Apptainer images with software versions optimized for GPU-accelerated Machine Learning frameworks. Users can clone these environments and add project-specific dependencies, typically related to data loading and visualization, and independent of the GPU software stack.A notable exception is represented by the software environment to develop Quantum Machine Learning (QML), featuring Python modules that simulate the effect of quantum operators on GPU and therefore requiring the same attention as other GPU-accelerated ML libraries to match the versions of the underlying software.In addition, Apptainer images specialized for the data processing of the LHC experiments can be obtained via CVMFS.
@@
A dedicated monitoring and accounting system has been set up for the platform in order to effectively control the use of all the platform’s resources and in particular of the GPUs. Several metric exporters have been configured to collect the information of interest and then expose it to a Prometheus Prometheus-link instance running in the platform. Some of these exporters are already available as Free and Open Source Software, such as Kube Eagle KubeEagle , which manages information about the use of the cluster’s CPUs and memory resources by the various components of $\mathrm{Kubernetes}$, or NVIDIA GPU DCGM exporter DCGM . Other exporters were developed on purpose, for example to monitor the usage of storage resources. All the metrics collected by Prometheus are then made visible and accessible through a Grafana grafana-docs dashboard. Grafana is run in a VM independent of the platform cluster and is used to monitor other VMs in the $\mathrm{AI\_INFN}$ OpenStack tenancy. It also hosts a PostgreSQL database for the accounting metrics, updated at regular intervals by averaging the metrics obtained from the monitoring Prometheus service.
@@ 4 Offloading: scale the applications beyond cluster boundaries
Users are allowed to scale beyond their notebook instance by creating $\mathrm{Kubernetes}$ jobs, enqueued and assigned to either local or remote resources by the Kueue controller Kueue . Kueue is designed to use local resources in an opportunistic way, configuring the running batch jobs to be immediately evicted in case new notebook instances are spawned pushing the cluster in a condition of resource contention. User do not create jobs directly accessing $\mathrm{Kubernetes}$ APIs, but passing through a dedicated microservice, named vkd, that validates user’s request based on memberships criteria and manage $\mathrm{Kubernetes}$ secrets that are not intended to be exposed to users, but still are needed for their jobs to be executed in the platform.
@@
Figure 1: The $\mathrm{AI\_INFN}$ platform enables offloading using InterLink as a virtual Kubelet provider to offload workloads from a Kubernetes cluster to external resources. InterLink interfaces with HTCondor, Slurm, and Podman via dedicated plugins, providing transparent access to HPC and cloud infrastructures.
@@
Virtual nodes are $\mathrm{Kubernetes}$ nodes that are not backed by a Linux kernel but mimic a $\mathrm{Kubernetes}$ kubelet in the interactions with the $\mathrm{Kubernetes}$ API server. The software component providing this interface is named Virtual Kubelet vk and is designed to ease the integration with various resource providers. The $\mathrm{AI\_INFN}$ platform relies on the InterLink Ciangottini:chep2024 provider. A further abstraction layer defining a simplified set of REST APIs that can be implemented by the so-called InterLink plugins providing the actual access to the compute resources. At the time of writing, the $\mathrm{AI\_INFN}$ platform is interfaced with plugins accessing HTCondor htcondor , Slurm slurm and Podman podman resources. Following a recent integration test, a $\mathrm{Kubernetes}$ plugin will be brought to production soon. A schematic representation of the architecture is provided in Figure 2.
@@
Figure 2 reports a recent scalability test involving resources provisioned by four different sites, without distributing the file system and for CPU-only payloads of the LHCb Flash Simulation Barbetti:2906203 . The plot illustrates the increase in job counts and their distribution across multiple computing sites. Stable contributions were maintained throughout the test by the INFN-Tier1 at CNAF in Bologna, provisioned via HTCondor. Comparable stability was observed from a Podman-based VM, whereas CINECAâs Leonardo supercomputer, provisioned via Slurm, showed greater variability in job activity during the initial hours. Later in the test, the Terabit HPC-Bubble, also provisioned via Slurm, was integrated and demonstrated consistent, sustained usage. The pronounced increase in job activity during the evening suggests improved scheduling efficiency. Overall, the test confirms the effectiveness of integrating multiple sites with heterogeneous provisioning mechanisms.